# RAISE: Enhancing Scientific Reasoning in LLMs via Step-by-Step Retrieval

### v2

@@ 4.2 Main Results
To assess the generalizability of RAISE across different LLM scales, we evaluate its performance on GPQA using LLaMA 3.1-8B Grattafiori et al. (2024) and GPT-4o mini OpenAI (2024), in addition to Mistral (used in our main experiments). As shown in Table 2, RAISE demonstrates consistent improvements over other baselines, exhibiting a similar trend to our main results with Mistral-24B. This shows that the effect of RAISE is not limited to a specific type of LLM, but can be applied to various LLMs with different scales.
@@
What is the concentration of OH ions in a solution of 0.3 M Ba(OH)2? Normality is an ambiguous measure of the concentration of a solution. It needs a definition of the equivalence factor… Barium hydroxide is a chemical compound with the formula Ba(OH)2(H2O). Barium hydroxide can be prepared by dissolving BaO in water… The Ba centers adopt a square anti-prismatic geometry. The RAISE-retrieved document clearly identifies barium hydroxide as $\mathrm{Ba(OH)_{2}}$ and explains its dissociation behavior in water, directly supporting the calculation of $[\mathrm{OH}^{-}]$ concentration.
Table 3: Examples where RAISE-retrieved documents provide logically relevant information for scientific reasoning compared to baseline RAG retrieval.
4.3 Analysis of RAISE
4.3.1 Impact of Problem Decomposition.
4.3.2 Qualitative Evaluation of Retrieved Documents
We qualitatively demonstrate the examples when RAISE retrieves logically relevant documents compared to convential RAG in Table 3. While RAG often retrieves documents that are topically related yet fail to address the reasoning needs of the subquestion, RAISE consistently identifies documents that include essential scientific principles, mechanisms, or equations. For instance, in questions involving chemical reactions, RAISE surfaces materials that explain the specific reactivity or the retarded time calculation, whereas RAG may return vague definitions or unrelated economic concepts. These cases illustrate how RAISE’s retrieval is not only domain-aware but also aligned with the logical demands of solving complex scientific problems.
@@ 4.3.2 Qualitative Evaluation of Retrieved Documents
Additionally, since LLM-as-judge may struggle with assessing logical relevance in expert-level science problems like GPQA and SuperGPQA, we conducted human evaluations with 2–3 experts per question, including chemistry professors and PhDs.The results can be found in Appendix D.2. These human evaluation results further confirm that RAISE significantly outperforms other baselines in selecting logically relevant documents.
5 Related Works
@@
Recent works have shown that LLMs can be applied for challenging scientific reasoning tasks. Unlike other domains, scientific reasoning requires not only step-by-step thinking, but also knowledge of specialized terminology and adaptation to continually evolving knowledge. Due to this challenging nature, many benchmarks have been proposed recently to tackle scientific reasoning with LLMs Rein et al. (2024); Zhong et al. (2025); Zhang et al. (2025b); Wang et al. (2024a); Jiang et al. (2024). Many works enhance scientific reasoning capabilities of LLMs through domain-specific training Taylor et al. (2022); Prabhakar et al. (2025); Zhang et al. (2024b), step-by-step reasoning Rueda et al. (2025); Rein et al. (2024), or retrieval of external knowledge or tools Ma et al. (2024); Zhong et al. (2025); Wellawatte et al. (2025); Li et al. (2025). Unlike previous works, we focus on applying step-by-step document retrieval from in-the-wild corpus without assuming access to well-curated and domain-specific corpus.
@@
A growing body of research has shown that decomposing complex problems into structured intermediate steps can enhance the reasoning abilities of LLMs. An influential early approach, Chain-of-Thought prompting Wei et al. (2022), introduced explicit, sequential reasoning steps, making the model’s thought process more transparent and coherent. This inspired methods such as Plan-and-Solve Wang et al. (2023), which emphasizes high-level planning before answering, and Step-Back Prompting Zheng et al. (2024), which encourages abstraction by prompting the model to reflect before solving. Least-to-Most prompting Zhou et al. (2023) extends this by breaking down tasks into simpler subproblems, solved in increasing order of difficulty.
@@
Retrieval-Augmented Generation (RAG) Lewis et al. (2020b) was initially proposed to improve LLMs’ factual accuracy and knowledge by retrieving relevant external documents during generation Lewis et al. (2020b); Karpukhin et al. (2020); Izacard and Grave (2021); Khattab and Zaharia (2020).
Recently, RAG has been extended for multi-hop reasoning, performing retrieval iteratively at multiple reasoning steps Liu et al. (2025); Zhao et al. (2024b); Zheng et al. (2024). In parallel, query reformulation and expansion techniques have been developed to enhance retrieval. Instead of using the original question, models generate enriched queries through prompting, such as intermediate answers or summaries. For example, HyDE Gao et al. (2023a) and CSQE Lei et al. (2024) demonstrate that carefully crafted queries greatly improve retrieval in complex, multi-step tasks.
@@
6 Conclusion and Limitations
We introduce RAISE, a step-by-step retrieval framework for scientific reasoning. We first decompose the problem into multiple subquestions and search queries, and then generate logical queries and retrieve logically relevant documents from in-the-wild corpus. We demonstrate the effectiveness of RAISE on three scientific reasoning benchmarks by comparing with various baselines. Our analysis shows that RAISE retrieves documents that are not only relevant in terms of the domain (e.g. definition of specificalized terminology) but also logically relevant documents for each subquestion, assisting the step-by-step reasoning process required for scientific reasoning.
However, RAISE has some limitations and areas for future improvement. First, we used only DPR for retrieval. Incorporating recent models Gao et al. (2023b); Basu et al. (2024); Shi et al. (2023) may further improve performance. In addition, our study focuses on scientific reasoning. Applying RAISE to other domains such as legal, commonsense, or medical reasoning could help evaluate its generalizability. Moreover, RAISE requires multiple LLM calls during logical query generation. Reducing this cost through better prompting, caching, or lightweight planning could improve scalability and real-world usability.
Appendix A RAISE Algorithm
Output: Final answer $y$
Step 1: Problem Decomposition
Generate subquestions and initial queries:
$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1})$
for $i=1$ to $n$ do
 Step 2: Logical Query Generation
   Reformulate initial query:
    $q_{i}^{*}\sim P_{\theta}(\cdot\mid r_{i},q_{i},p_{2})$
 Step 3: Knowledge Retrieval
[…]
    $D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$
 Step 4: Subquestion Answering
 if $i=1$ then
       $a_{i}\sim P_{\theta}(\cdot\mid x,r_{1},D_{1},p_{3})$
 else
       $a_{i}\sim P_{\theta}(\cdot\mid x,\{(r_{j},a_{j})\}_{j=1}^{i-1},r_{i},D_{i},p_{3})$
   end if
end for
Step 5: Final Answer Composition
Generate final answer using all subanswers:
$y\sim P_{\theta}(\cdot\mid x,\{(r_{i},a_{i})\}_{i=1}^{n},p_{4})$
Algorithm 1 RAISE Inference Procedure
Appendix B Dataset Details
GPQA
Rein et al. (2024) This dataset consists of physics, biology, and chemistry questions written by domain experts. We use GPQA diamond subset, which consist of 198 high-quality questions selected based on human performance. Specifically, this subset includes questions that both experts answer correctly while the majority of non-experts fail to solve. Each question typically demands multi-step reasoning, precise formula manipulation, and access to external scientific facts (e.g., physical constants, definitions). Due to its alignment with our target setting, GPQA serves as the primary evaluation benchmark throughout our experiments.
SuperGPQA
Team et al. (2025) SuperGPQA is a large-scale benchmark designed to evaluate graduate-level reasoning across 13 disciplines, 72 fields, and 285 graduate-level disciplines. In alignment with the scientific reasoning focus of our work, we select science and engineering domains for evaluation. Each domain is further divided by three difficulty levels(easy, medium, and hard). To reduce computational overhead while maintaining consistency, we randomly sample 199 questions per subset using a fixed seed (42). Specifically, our experiments include 199 examples each from science-hard, science-middle, and engineering-hard subsets.
MMLU
Hendrycks et al. (2021); Wang et al. (2024b) The MMLU benchmark covers a wide range of subjects across multiple domains. For out experiments, we focus on STEM and Professional categories. The STEM contains university-level science and engineering subjects such as college mathematics and computer science, while the Professional category covers specialized fields that typically require professional training or advanced education, including law, medicine, and chemistry. We specifically select three subsets: college chemistry from MMLU-STEM and professional chemistry and biology from MMLU-Pro. These subsets are chosen to evaluate our method’s ability to perform scientific reasoning in both academic and professional contexts involving complex domain knowledge.
Appendix C Experiment Details
C.1 Baseline Details
CoT Wei et al. (2022); Kojima et al. (2023) We apply Chain-of-Thought prompting for direct reasoning, where the model is encouraged to explicitly generate intermediate reasoning steps through prompting (Think step by step).
CoT+RAG
Lewis et al. (2020b) We implement CoT+RAG by combining Chain-of-Thought prompting with retrieval, where the model is prompted to solve the problem step-by-step while also leveraging external knowledge. Specifically, we provide the model with a CoT-style prompt encouraging step-by-step reasoning, alongside the original question and documents retrieved using the original question as the search query.
Least-to-Most
Zhou et al. (2023) Least-to-Most is a decomposed reasoning strategy that breaks down a complex problem into a sequence of simpler subquestions, which are then solved sequentially without retrieval augmentation. This subquestion decomposition pipeline serves as the foundational structure for other decomposed reasoning methods as well.
Step-Back
(Decomposed reasoning) Zheng et al. (2024) We implement Step-Back for decomposed reasoning by applying the Step-Back prompting method to each subquestion in a decomposed reasoning framework. While the original Step-Back paper does not cover the application of this method to decomposed subquestions, we extend it for a fair comparison with our approach. Specifically, after decomposing the original question into subquestions, we use the Step-Back prompting strategy to extract a high-level principle for each subquestion, and then provide the subquestion along with its corresponding principle to guide the model’s reasoning.
Least-to-Most+RAG
(Decomposed reasoning with RAG) Liu et al. (2025) We implement RAG by first decomposing the original problem into subquestions and then retrieving documents using each subquestion as a query. The retrieved documents are provided to the model along with the corresponding subquestion to support its reasoning.
Step-Back+RAG
(Decomposed reasoning with RAG) Zheng et al. (2024) We extend the Step-Back prompting strategy to a retrieval-augmented setting for fair comparison with our method. After decomposing the original question into subquestions, we generate a principle abstraction for each subquestion using Step-Back prompting, and use it as a query to retrieve evidence. The retrieved documents are then provided alongside the original subquestion to guide the model’s reasoning.
HyDE
(Decomposed reasoning with RAG) Gao et al. (2023a) We apply the HyDE approach to each subquestion in a decomposed reasoning framework. For each subquestion, the model first generates a hypothetical answer, which is then used as a query to retrieve supporting documents. The retrieved evidence, together with the subquestion, is provided to the model to support step-by-step reasoning.
C.2 Retriever Configuration
We use the pre-trained DPR encoder from the ’facebook/dpr-question_encoder-single-nq-base’ model Karpukhin et al. (2020), which is a BERT-based encoder trained for open-domain question answering. This encoder is trained on the Natural Question (NQ) dataset Kwiatkowski et al. (2019) and is designed to map questions into 768-dimensional dense vector representations for retrieval.
For the retrieval corpus, we use the preprocessed Wikipedia passages provided by ’facebook/wiki_dpr’ Karpukhin et al. (2020), a corpus widely used to evaluate DPR-based retrieval models. This corpus is constructed from the December 20, 2018 Wikipedia dump, where each article is split into multiple, disjoint text blocks of 100 words, resulting in approximately 21 million passages. Each passage is accompanied by the title of the wikipedia page it comes from along with DPR embedding.
To enable efficient retrieval over the passage embeddings, we use an exact FAISS index. FAISS (Facebook AI Similarity Search) Johnson et al. (2019); Douze et al. (2024) is widely used library for fast similarity search over dense vectors.
Appendix D Additional Results
D.1 Further Analysis on GPQA
Figure 5: Domain-wise accuracy comparison between CoT and RAISE on the GPQA Diamond subset
D.2 Human Evaluation of Retrieved Documents
Previously, in Section 4.3.2, we evaluated the logical relevancy of documents retrieved by the baseline methods and RAISE using the LLM-as-a-judge approach. However, since our domain includes complex, expert-level questions, and LLM-based evaluations may overlook domain-specific reasoning and often rely on surface-level features, we supplemented our analysis with a small-scale human evaluation of 20 subquestion–document pairs. Each pair was assessed by at least three annotators, including Ph.D. students and a faculty member in chemistry, with the method provenance concealed to maintain objectivity. The results are presented in Figure 6.
As also discussed in the LLM-as-a-judge results, the human evaluation indicates that RAISE produces significantly fewer irrelevant documents compared to all other methods, while achieving the highest proportion of logically relevant documents. Although limited in scale due to time and cost constraints, we believe this evaluation provides meaningful human validation of RAISE’s effectiveness and serves as a valuable complement to the LLM-based assessments.

### v3

@@ Appendix C Additional Results
To assess the generalizability of RAISE across different LLM scales, we evaluate its performance on GPQA using LLaMA 3.1-8B [4] and GPT-4o mini [26], in addition to Mistral (used in our main experiments). As shown in Table 2, RAISE demonstrates consistent improvements over other baselines, exhibiting a similar trend to our main results with Mistral-24B. This shows that the effect of RAISE is not limited to a specific type of LLM, but can be applied to various LLMs with different scales.
@@ C.2 Qualitative Evaluation of Retrieved Documents
What is the concentration of OH ions in a solution of 0.3 M Ba(OH)2? Normality is an ambiguous measure of the concentration of a solution. It needs a definition of the equivalence factor… Barium hydroxide is a chemical compound with the formula Ba(OH)2(H2O). Barium hydroxide can be prepared by dissolving BaO in water… The Ba centers adopt a square anti-prismatic geometry. The RAISE-retrieved document clearly identifies barium hydroxide as $\mathrm{Ba(OH)_{2}}$ and explains its dissociation behavior in water, directly supporting the calculation of $\mathrm{OH}^{-}]$ concentration.
Figure 3: Examples where RAISE-retrieved documents provide logically relevant information for scientific reasoning compared to baseline RAG retrieval.
We qualitatively demonstrate the examples when RAISE retrieves logically relevant documents compared to convential RAG in Figure 3. While RAG often retrieves documents that are topically related yet fail to address the reasoning needs of the subquestion, RAISE consistently identifies documents that include essential scientific principles, mechanisms, or equations. For instance, in questions involving chemical reactions, RAISE surfaces materials that explain the specific reactivity or the retarded time calculation, whereas RAG may return vague definitions or unrelated economic concepts. These cases illustrate how RAISE’s retrieval is not only domain-aware but also aligned with the logical demands of solving complex scientific problems.
Figure 4: Examples comparing query generation methods (Step-Back+RAG, HyDE, and RAISE) for the same subquestion. Both Step-Back+RAG and HyDE are methods that reformulate the original query to retrieve more relevant documents. These methods are included as baselines in the main comparison table.
@@ C.3 Evaluation of Logical Relevancy of Retrieved Documents
Since our domain includes complex, expert-level questions, and LLM-based evaluations may overlook domain-specific reasoning and often rely on surface-level features, we supplemented our analysis with a small-scale human evaluation of 20 subquestion–document pairs. Each pair was assessed by at least three annotators, including Ph.D. students and a faculty member in chemistry, with the method provenance concealed to maintain objectivity. As also discussed in the LLM-as-a-judge results, the human evaluation indicates that RAISE produces significantly fewer irrelevant documents compared to all other methods, while achieving the highest proportion of logically relevant documents. Although limited in scale due to time and cost constraints, we believe this evaluation provides meaningful human validation of RAISE’s effectiveness and serves as a valuable complement to the LLM-based assessments.
C.4 Further Analysis on GPQA
Figure 7: Domain-wise accuracy comparison between CoT and RAISE on the GPQA Diamond subset.
Appendix D Related Works
@@
Recent works have shown that LLMs can be applied for challenging scientific reasoning tasks. Unlike other domains, scientific reasoning requires not only step-by-step thinking, but also knowledge of specialized terminology and adaptation to continually evolving knowledge. Due to this challenging nature, many benchmarks have been proposed recently to tackle scientific reasoning with LLMs [29, 50, 46, 35, 10]. Many works enhance scientific reasoning capabilities of LLMs through domain-specific training [31, 27, 43], step-by-step reasoning [30, 29], or retrieval of external knowledge or tools [24, 50, 38, 22]. Unlike previous works, we focus on applying step-by-step document retrieval from in-the-wild corpus without assuming access to well-curated and domain-specific corpus.
@@
A growing body of research has shown that decomposing complex problems into structured intermediate steps can enhance the reasoning abilities of LLMs. An influential early approach, Chain-of-Thought prompting [37], introduced explicit, sequential reasoning steps, making the model’s thought process more transparent and coherent. This inspired methods such as Plan-and-Solve [33], which emphasizes high-level planning before answering, and Step-Back Prompting [49], which encourages abstraction by prompting the model to reflect before solving. Least-to-Most prompting [51] extends this by breaking down tasks into simpler subproblems, solved in increasing order of difficulty.
@@
Retrieval-Augmented Generation (RAG) [21] was initially proposed to improve LLMs’ factual accuracy and knowledge by retrieving relevant external documents during generation [21, 14, 8, 15].
Recently, RAG has been extended for multi-hop reasoning, performing retrieval iteratively at multiple reasoning steps [23, 48, 49]. In parallel, query reformulation and expansion techniques have been developed to enhance retrieval. Instead of using the original question, models generate enriched queries through prompting, such as intermediate answers or summaries. For example, HyDE [3] and CSQE [19] demonstrate that carefully crafted queries greatly improve retrieval in complex, multi-step tasks.